Hello Tutors. Can I get some help with this homework problem please.
Given that P(A) = 0.6 and P(B) = 0.3, and P(A and B) = 0.18, determine P(B|A

Answer by stanbon(75887)

Given that P(A) = 0.6 and P(B) = 0.3, and P(A and B) = 0.18, determine P(B|A)
-----------------
P(B | A) = P(B and A)/P(A) = 0.18/0.6 = 0.3
===================
Cheers,
Stan H.
================